OpenStudy (anonymous):

c(c+2)=0 solve using the priciple of zero products use comma to seperate answers as needed. type each solution only once.

OpenStudy (anonymous):

\[c \in(0,-2)\]

myininaya (myininaya):

you have c(c+2)=0 set both factors =0 c=0 , c+2=0

OpenStudy (anonymous):

Im confused!!!

OpenStudy (anonymous):

ok like X(X-3)=0 the only way this will work is if x=0 or x=3 0(0-3)=0 check and 3(3-3)=0 check any factor outside the ( ) along is almost always 0 the take the opposite of whats inside the ( )